Find 2 numbers such that their sums are77 and their difference is 25
Answer by Abu peterson ayuba(1)

Let the two numbers be x and y. x+y=77--(1) and x-y=25--(2) > x=77-y--(3) then substituate equation 3 into equation 2. (77-y)-y=25 >77-2y=25 >-2y=25-77 >-2y=-52 (then divid both sides by -2)> y=26. To get x,substitute y=26 into equation 1 >x+y=77 >x+26=77 >x=77-26 >x=51. (x=51 and y=26)